# Integral means spectrum functionals on Teichmüller spaces

### v11

@@ 1. Introduction
Let $\Delta=\{z:|z|<1\}$ denote the unit disk in the complex plane $\mathbb{C}$. We denote the extended complex plane by $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$. For two positive numbers $A,B$, we write $A\asymp B$ if there are two constants $C_{1}>0,C_{2}>0$ which are independent on the arguments and such that $C_{1}B\leq A\leq C_{2}B$.
We denote the class of all univalent functions $f$ in $\Delta$ by $\mathcal{U}$. Let $\mathcal{S}$ be the class of all univalent functions $f$ in $\Delta$ with $f(0)=f^{\prime}(0)-1=0$. We let $\mathcal{S}_{b}$ be the subclass of $\mathcal{S}$ which consists of all bounded univalent functions.
@@
The universal integral mean spectrums $B(t)$ and $B_{b}(t)$ are defined as
@@
It is an important problem in the theory of univalent functions(conformal mappings) to determine the exact values of the universal integral mean spectrums $B(t)$ and $B_{b}(t)$. We review some known results and open problems on this topic. It was proved by Makarov in [29] that
@@
We let $\mathcal{S}_{q}$ be the class of all univalent functions $f$ that belong to $\mathcal{S}_{b}$ and admit a quasiconformal extension to $\widehat{\mathbb{C}}$. Later, if a univalent function $f$ belongs to $\mathcal{S}_{q}$, we still use $f$ to denote its quasiconformal extension. Recently, some researchers have studied the integral means spectrums of univalent functions admitting a quasiconformal extension to $\widehat{\mathbb{C}}$, see for example, [20], [21], [22], [23], [24], [35] and [36]. By the fractal approximation principle, see [29], [8], we know that
@@
(1) For fixed $t\in\mathbb{R}$ with $t\neq 0$, how does the integral means spectrum $\beta_{f}(t)$ depend on $f$ in the class $\mathcal{S}_{q}$? (2) For each $t\neq 0$, does there exist at least one extremal function for $B_{b}(t)$? (3) If the extremal functions for $B_{b}(t)$ exist, which subset of $S$ do they belong to?
@@ 2.1. Univalent functions (conformal mappings)
For two locally univalent function $f_{1},f_{2}$ in $\Delta$, we have $N_{f_{1}}(z)=N_{f_{2}}(z)$ for all $z\in\Delta$ if and only if there is two number $a,b\in\mathbb{C}$ such that $f_{1}(z)=af_{2}(z)+b$ for all $z\in\Delta$.
The if part of this claim is easy. We only prove the only if part. We let $N_{f_{j}}(z):=\phi(z),j=1,2$. Then
@@
for each $z\in\Delta$. We will show that there is an $f_{\Phi}\in\mathcal{U}$ such that $N_{f_{\Phi}}=\Phi$. Without loss of generality, since the Pre-Schwarzian derivative is affine invariant, we assume that ${\bf f}_{n}\in\mathcal{S}$ for all $n$. Noting that $\{{\bf f}_{n}\}$ is a normal family, then $\{{\bf f}_{n}\}$ contains a subsequences (still denoted by $\{{\bf f}_{n}\}$) which is locally uniformly convergence in $\Delta$. We denote by $g$ the limit of the subsequences, i.e.,
@@ 2.2. Quasiconformal mappings
We say a sense-preserving homeomorphism $f$, from one open domain $\Omega$ in $\mathbb{C}$ to another, is a quasiconformal mapping if it has locally square integral distributional derivatives and satisfies the Beltrami equation $\bar{\partial}f=\mu_{f}\partial{f}$ with
@@
We call $D_{f}(z)$ the dilatation function of $f$ in $\Omega$ and $K_{f}$ is called as the dilatation of $f$.
@@
It is easy to see that $2^{\mathbf{m}+1}\mathbf{d}_{0}\geq\frac{1}{2}\mathbf{d}$. By the Pompieu’s formula, we have
@@
Here $J_{\mathbf{h}}$ is the Jacobian of $\mathbf{h}$. It follows from Cauchy-Schwartz’s inequality that
@@
Since $\varepsilon<\frac{1}{3}(1-\|\mu_{f}\|_{\infty})-b(\mathbf{h})$ so that $b(\mathbf{h})+\varepsilon<\frac{1}{3}(1-\|\mu_{f}\|_{\infty})$. Then it follows from (2.3) that
@@
To continuous the proof, we need the following claim.
@@
for any two different points $z_{1},z_{2}$ in $\overline{\Delta}$. On the other hand, since $\pi_{1}\circ{\mathbf{h}}^{-1}\circ{\pi_{2}}^{-1}$ is a quasiconformal mapping from $\Delta$ to itself with fixing the origin. We see from Lemma 2.11 that
@@
$(1-|z|^{2})|f^{\prime}(z)|\asymp{\textup{dist}}(f(z),f(\mathbb{T})))\asymp|f(z)-f(e^{i\arg{z}})|,$
$(1-|z|^{2})|g^{\prime}(z)|\asymp{\textup{dist}}(g(z),g(\mathbb{T})))\asymp|g(z)-g(e^{i\arg{z}})|,$
@@ 3. Integral means spectrum functionals on Teichmüller spaces
The mapping $\Lambda_{2}$ is known as Bers mappings. We call $\Lambda_{1}$ Pre-Bers mapping. It is well known that
@@
It is obvious that $\Lambda_{1}$ is bijective. We will first show that $\Lambda_{1}$ is continuous. For $\mu,\nu\in\mathcal{M}(\Delta^{*})$. We may assume that $\|\mu-\nu\|_{\infty}>0$, and let $\mathbf{f}=f_{\nu}\circ f_{\mu}^{-1}$. Then we see that $\mathbf{f}(0)=\mathbf{f}^{\prime}(0)-1=0,\mathbf{f}(\infty)=\mathbf{f}(\infty)$ and from (2.3) that
@@
We see from [31] that $|\Phi(t)|\leq 8$. Noting that $f(z,0)=z$ so that $\Phi(0)=0$, then by using Schwarz’s lemma, we obtain that $|\Phi(t)|\leq 8t\|\mu_{\mathbf{f}}\|_{\infty}$ for $t\in\mathcal{D}_{\mathbf{f}}$. On the other hand, from $f(z,1)=\mathbf{f}(z)$, we have $|N_{\mathbf{f}}(z)|\rho_{f_{\mu}(\Delta)}^{-1}(z)\leq 8\|\mu_{\mathbf{f}}\|_{\infty}$ for all $z\in f_{\mu}(\Delta)$. It follows from the fact $\mathbf{f}=f_{\nu}\circ f_{\mu}^{-1}$ and (2.1) that
@@
In view of Propposition 3.2 and 3.3, we can identify the universal Teichmüller space with $T_{1}$ or $T_{2}$. Let $\mathcal{S}_{q}^{\infty}$ be the class of all functions $f\in\mathcal{S}_{q}$ that have a quasiconformal extension (still denoted by $f$) to $\widehat{\mathbb{C}}$ with $f(\infty)=\infty$. We set
@@
We next identify the universal Teichmüller space with $T_{1}=\mathbf{N}_{q}$. From Proposition 2.1, since $\mathbf{N}$ is closed in $E_{1}$, then the closure $\overline{\mathbf{N}}_{q}$ of $\mathbf{N}_{q}$ is contained in $\mathbf{N}$. We will see $\overline{T}_{1}=\overline{\mathbf{N}}_{q}$ as a model of the closure of the universal Teichmüller space. For any $\phi\in\overline{\mathbf{N}}_{q}$, there is a unique univalent function $f_{\phi}(z)$ with $f_{\phi}\in\mathcal{S}$ and such that
@@
We continuous to prove the only if part. When $\mu$ is asymptotically equivalent to $\nu$, we know that there is a $\widetilde{\nu}\in\mathcal{M}(\Delta^{*})$ such that $\widetilde{\nu}\sim\nu$ and $\widetilde{\nu}(z)-\mu(z)\rightarrow 0$ as $z\rightarrow 1^{-}$. Let $\widetilde{\mathbf{h}}=f_{\widetilde{\nu}}\circ f_{\mu}^{-1}$. It follows from (2.3) that
@@
In the standard theory of universal asymptotic Teichmüller space, $AT$ is embedding mapped to an open subset of a complex Banach space by using the Bers mapping induced by the Schwarzian derivative. We shall consider the mapping induced by the Pre-Schwarzian derivative. We let
@@
The mapping $\widetilde{\Lambda}_{2}$ is called as asymptotic Bers map. It was proved in [14] that
@@
The mapping $\widetilde{\Lambda}_{1}$ is called as asymptotic Pre-Bers mapping.
@@
Noting that $d_{AT}([\mu]_{AT},[\nu]_{AT})\leq d_{T}([\mu]_{T},[\nu]_{T})$ for any $\mu,\nu\in\mathcal{M}(\Delta^{*})$, we see that the statement that $I_{AT}$ is continuous on $AT$ also implies Theorem 3.6.
@@ 5.1. A final main theorem
Let $\mu\in\mathcal{M}(\Delta^{*})$ and let $k=\|\mu\|_{\infty}$. Then there exists a (canonical) holomorphic motion $\mathbf{H}:\Delta\times\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ such that $\mathbf{H}(k,z)=f_{\mu}(z)$. Moreover, when $k=\|\mu\|_{\infty}>0$, for each fixed $\lambda\in\Delta$, $\mathbf{H}(\lambda,z)$ is a qusiconformal mapping from $\widehat{\mathbb{C}}$ to itself with $\mu_{\mathbf{H}(\lambda,z)|_{\Delta}}=0,\,\,z\in\Delta$ and $\mu_{\mathbf{H}(\lambda,z)|_{\Delta^{*}}}=\lambda\cdot\frac{\mu(z)}{\|\mu\|_{\infty}},\,z\in\Delta^{*}.$
@@
is subharmonic in $\Delta$. Noting that
@@
Case 1. If $I_{\mathcal{E}}\geq\frac{1}{2}\mathcal{A}_{n}$, then, for $\Delta t>0$, we have
@@
This corollary tells us that the extremal function for $B_{b}(t)$ can not be from the class $\mathcal{S}_{q}$ when $t\neq 0$. This partially answer part (3) of Question 1.13.
@@ 5.2. Remarks
We finally identify the universal Teichmüller space with $T_{2}=\mathbf{S}_{q}$. Hence we can see $\overline{T}_{2}=\overline{\mathbf{S}}_{q}$ as another model of the closure of the universal Teichmüller space. From [3], we know that the boundary $\partial\mathbf{S}_{q}$ of $\mathbf{S}_{q}$ is larger than the one of $\mathbf{N}_{q}$. If $\phi\in\mathbf{S}_{q}$, then there is a unique univalent function $f_{\phi}$ with $f_{\phi}\in\mathcal{S}_{q}^{\infty}$ and such that $\phi(z)=S_{f_{\phi}}(z).$ Hence, the IMS functional $I_{T_{2}}:\phi\mapsto\beta_{f_{\phi}}(t)$ is well-defined on $\mathbf{S}_{q}$. From Proposition 3.2, 3.3 and Theorem 3.6, we obtain that
@@
If $\phi\in\partial\mathbf{S}_{q}$, we know that there is a sequence $\{f_{n}(z)\}_{n=1}^{\infty}$, $f_{n}\in\mathcal{S}_{q}^{\infty}$ such that $\lim_{n\rightarrow\infty}\|S_{f_{n}}-\phi\|_{E_{2}}=0,$ and for each $z\in\Delta$, the sequence $\{f_{n}(z)\}_{n=1}^{\infty}$ converges. Then, take $f_{\phi}(z)=\lim_{n\rightarrow\infty}f_{n}(z),z\in\Delta$, we see that $f_{\phi}\in\mathcal{S}$ with $\phi(z)=S_{f_{\phi}}(z)$. In view of the normalization, we believe that $f_{\phi}$ should be unique, here the statement $f_{\phi}$ is unique means that, if there is another sequence $\{\widehat{f}_{n}(z)\}_{n=1}^{\infty}$, $\widehat{f}_{n}\in\mathcal{S}_{q}^{\infty}$ such that $\lim_{n\rightarrow\infty}\|S_{\widehat{f}_{n}}-\phi\|_{E_{2}}=0,$ and for each $z\in\Delta$, the sequence $\{\widehat{f}_{n}(z)\}_{n=1}^{\infty}$ converges, then, take $\widehat{f}_{\phi}(z)=\lim_{n\rightarrow\infty}\widehat{f}_{n}(z)$, we have $\widehat{f}_{\phi}(z)=f_{\phi}(z)$ for any $z\in\Delta$. But we have not found a proof for this statement. We leave it as the following conjecture, which seems to be non-trivial.

### v12

@@ 1. Introduction
Let $\Delta=\{z:|z|<1\}$ denote the unit disk in the complex plane $\mathbb{C}$. We denote the extended complex plane by $\widehat{\mathbb{C}}$$\mathbb{C}\cup\{\infty\}$. For two positive numbers $A,B$, we write $A\asymp B$ if there are two constants $C_{1}>0,C_{2}>0$ which are independent of the arguments and such that $C_{1}B\leq A\leq C_{2}B$.
We denote the class of all univalent functions $f$ in $\Delta$ by $\mathcal{U}$. Let $\mathcal{S}$ be the class of all univalent functions $f$ in $\Delta$ with $f(0)=f^{\prime}(0)-1=0$. We let $\mathcal{S}_{b}$ be the subclass of $\mathcal{S}$, which consists of all bounded univalent functions.
@@
The universal integral mean spectrum $B(t)$ and $B_{b}(t)$ are defined as
@@
It is an important problem in the theory of univalent functions(conformal mappings) to determine the exact values of the universal integral mean spectra $B(t)$ and $B_{b}(t)$. We review some known results and open problems on this topic. It was proved by Makarov in [29] that
@@
We let $\mathcal{S}_{q}$ be the class of all univalent functions $f$ that belong to $\mathcal{S}_{b}$ and admit a quasiconformal extension to $\widehat{\mathbb{C}}$. Later, if a univalent function $f$ belongs to $\mathcal{S}_{q}$, we still use $f$ to denote its quasiconformal extension. Recently, some researchers have studied the integral means spectra of univalent functions admitting a quasiconformal extension to $\widehat{\mathbb{C}}$, see for example, [20], [21], [22], [23], [24], [35] and [36]. By the fractal approximation principle, see [29], [8], we know that
@@
(1) For fixed $t\in\mathbb{R}$ with $t\neq 0$, how does the integral means spectrum $\beta_{f}(t)$ depend on $f$ in the class $\mathcal{S}_{q}$? (2) For each $t\neq 0$, does there exist at least one extremal function for $B_{b}(t)$? (3) If the extremal functions for $B_{b}(t)$ exist, which subset of $\mathcal{S}$ do they belong to?
@@ 2.1. Univalent functions (conformal mappings)
For two locally univalent functions $f_{1},f_{2}$ in $\Delta$, we have $N_{f_{1}}(z)=N_{f_{2}}(z)$ for all $z\in\Delta$ if and only if there are two numbers $a,b\in\mathbb{C}$ such that $f_{1}(z)=af_{2}(z)+b$ for all $z\in\Delta$.
The if part of this claim is trivial. We only prove the only if part. We let $N_{f_{j}}(z):=\phi(z),j=1,2$. Then
@@
for each $z\in\Delta$. We will show that there is an $f_{\Phi}\in\mathcal{U}$ such that $N_{f_{\Phi}}=\Phi$. Without loss of generality, since the Pre-Schwarzian derivative is affine invariant, we assume that ${\bf f}_{n}\in\mathcal{S}$ for all $n$. Noting that $\{{\bf f}_{n}\}$ is a normal family, then $\{{\bf f}_{n}\}$ contains a subsequence (still denoted by $\{{\bf f}_{n}\}$) which is locally uniformly convergent in $\Delta$. We denote by $g$ the limit of the subsequence, i.e.,
@@ 2.2. Quasiconformal mappings
We say a sense-preserving homeomorphism $f$, from one open domain $\Omega$ in $\mathbb{C}$ to another is a quasiconformal mapping if it has locally square integral distributional derivatives and satisfies the Beltrami equation $\bar{\partial}f=\mu_{f}\partial{f}$ with
@@
We call $D_{f}(z)$ the dilatation function of $f$ in $\Omega$ and $K_{f}$ is called the dilatation of $f$.
@@
It is easy to see that $2^{\mathbf{m}+1}\mathbf{d}_{0}\geq\frac{1}{2}\mathbf{d}$. By Pompeiu’s formula, we have
@@
Here $J_{\mathbf{h}}$ is the Jacobian of $\mathbf{h}$. It follows from Cauchy-Schwarz’s inequality that
@@
Since $\varepsilon<\frac{1}{3}(1-\|\mu_{f}\|_{\infty})-b(\mathbf{h})$ we have $b(\mathbf{h})+\varepsilon<\frac{1}{3}(1-\|\mu_{f}\|_{\infty})$. Then it follows from (2.3) that
@@
To continue the proof, we need the following claim.
@@
for any two different points $z_{1},z_{2}$ in $\overline{\Delta}$. On the other hand, since $\pi_{1}\circ{\mathbf{h}}^{-1}\circ{\pi_{2}}^{-1}$ is a quasiconformal mapping from $\Delta$ to itself fixing the origin. We see from Lemma 2.11 that
@@
$(1-|z|^{2})|f^{\prime}(z)|\asymp{\textup{dist}}(f(z),f(\mathbb{T}))\asymp|f(z)-f(e^{i\arg{z}})|,$
$(1-|z|^{2})|g^{\prime}(z)|\asymp{\textup{dist}}(g(z),g(\mathbb{T}))\asymp|g(z)-g(e^{i\arg{z}})|,$
@@ 3. Integral means spectrum functionals on Teichmüller spaces
The mapping $\Lambda_{2}$ is known as the Bers mapping. We call $\Lambda_{1}$ the Pre-Bers mapping. It is well known that
@@
It is obvious that $\Lambda_{1}$ is bijective. We will first show that $\Lambda_{1}$ is continuous. For $\mu,\nu\in\mathcal{M}(\Delta^{*})$. We may assume that $\|\mu-\nu\|_{\infty}>0$, and let $\mathbf{f}=f_{\nu}\circ f_{\mu}^{-1}$. Then we see that $\mathbf{f}(0)=\mathbf{f}^{\prime}(0)-1=0,\mathbf{f}(\infty)=\infty$ and from (2.3) that
@@
We see from [31] that $|\Phi(t)|\leq 8$. Note that $f(z,0)=z$ so that $\Phi(0)=0$, then by using Schwarz’s lemma, we obtain that $|\Phi(t)|\leq 8t\|\mu_{\mathbf{f}}\|_{\infty}$ for $t\in\mathcal{D}_{\mathbf{f}}$. On the other hand, from $f(z,1)=\mathbf{f}(z)$, we have $|N_{\mathbf{f}}(z)|\rho_{f_{\mu}(\Delta)}^{-1}(z)\leq 8\|\mu_{\mathbf{f}}\|_{\infty}$ for all $z\in f_{\mu}(\Delta)$. It follows from the fact $\mathbf{f}=f_{\nu}\circ f_{\mu}^{-1}$ and (2.1) that
@@
In view of Proposition 3.2 and 3.3, we can identify the universal Teichmüller space with $T_{1}$ or $T_{2}$. Let $\mathcal{S}_{q}^{\infty}$ be the class of all functions $f\in\mathcal{S}_{q}$ that have a quasiconformal extension (still denoted by $f$) to $\widehat{\mathbb{C}}$ with $f(\infty)=\infty$. We set
@@
We next identify the universal Teichmüller space with $T_{1}=\mathbf{N}_{q}$. From Proposition 2.1, since $\mathbf{N}$ is closed in $E_{1}$, then the closure $\overline{\mathbf{N}}_{q}$ of $\mathbf{N}_{q}$ is contained in $\mathbf{N}$. We regard $\overline{T}_{1}=\overline{\mathbf{N}}_{q}$ as a model of the closure of the universal Teichmüller space. For any $\phi\in\overline{\mathbf{N}}_{q}$, there is a unique univalent function $f_{\phi}(z)$ with $f_{\phi}\in\mathcal{S}$ and such that
@@
We continue to prove the only if part. When $\mu$ is asymptotically equivalent to $\nu$, we know that there is a $\widetilde{\nu}\in\mathcal{M}(\Delta^{*})$ such that $\widetilde{\nu}\sim\nu$ and $\widetilde{\nu}(z)-\mu(z)\rightarrow 0$ as $z\rightarrow 1^{-}$. Let $\widetilde{\mathbf{h}}=f_{\widetilde{\nu}}\circ f_{\mu}^{-1}$. It follows from (2.3) that
@@
In the standard theory of universal asymptotic Teichmüller space, $AT$ is embedded into an open subset of a complex Banach space by using the Bers mapping induced by the Schwarzian derivative. We shall consider the mapping induced by the Pre-Schwarzian derivative. We let
@@
The mapping $\widetilde{\Lambda}_{2}$ is called asymptotic Bers mapping. It was proved in [14] that
@@
The mapping $\widetilde{\Lambda}_{1}$ is called asymptotic Pre-Bers mapping.
@@
Note that $d_{AT}([\mu]_{AT},[\nu]_{AT})\leq d_{T}([\mu]_{T},[\nu]_{T})$ for any $\mu,\nu\in\mathcal{M}(\Delta^{*})$, we see that the statement that $I_{AT}$ is continuous on $AT$ also implies Theorem 3.6.
@@ 5.1. A final main theorem
Let $\mu\in\mathcal{M}(\Delta^{*})$ and let $k=\|\mu\|_{\infty}$. Then there exists a (canonical) holomorphic motion $\mathbf{H}:\Delta\times\widehat{\mathbb{C}}\rightarrow\widehat{\mathbb{C}}$ such that $\mathbf{H}(k,z)=f_{\mu}(z)$. Moreover, when $k=\|\mu\|_{\infty}>0$, for each fixed $\lambda\in\Delta$, $\mathbf{H}(\lambda,z)$ is a quasiconformal mapping from $\widehat{\mathbb{C}}$ to itself with $\mu_{\mathbf{H}(\lambda,z)|_{\Delta}}=0,\,\,z\in\Delta$ and $\mu_{\mathbf{H}(\lambda,z)|_{\Delta^{*}}}=\lambda\cdot\frac{\mu(z)}{\|\mu\|_{\infty}},\,z\in\Delta^{*}.$
@@
is subharmonic in $\Delta$. Note that
@@
Case 1. If $I_{\mathcal{E}}\geq\frac{1}{2}\mathcal{A}_{n}$, then for $\Delta t>0$, we have
@@
This corollary tells us that the extremal function for $B_{b}(t)$ can not be from the class $\mathcal{S}_{q}$ when $t\neq 0$. This partially answers part (3) of Question 1.13.
@@ 5.2. Remarks
We finally identify the universal Teichmüller space with $T_{2}=\mathbf{S}_{q}$. Hence we can see $\overline{T}_{2}=\overline{\mathbf{S}}_{q}$ as another model of the closure of the universal Teichmüller space. From [3], we know that the boundary $\partial\mathbf{S}_{q}$ of $\mathbf{S}_{q}$ is larger than that of $\mathbf{N}_{q}$. If $\phi\in\mathbf{S}_{q}$, then there is a unique univalent function $f_{\phi}$ with $f_{\phi}\in\mathcal{S}_{q}^{\infty}$ and such that $\phi(z)=S_{f_{\phi}}(z).$ Hence, the IMS functional $I_{T_{2}}:\phi\mapsto\beta_{f_{\phi}}(t)$ is well-defined on $\mathbf{S}_{q}$. From Proposition 3.2, 3.3 and Theorem 3.6, we obtain that
@@
If $\phi\in\partial\mathbf{S}_{q}$, we know that there is a sequence $\{f_{n}(z)\}_{n=1}^{\infty}$, $f_{n}\in\mathcal{S}_{q}^{\infty}$ such that $\lim_{n\rightarrow\infty}\|S_{f_{n}}-\phi\|_{E_{2}}=0,$ and for each $z\in\Delta$, the sequence $\{f_{n}(z)\}_{n=1}^{\infty}$ converges. Then, take $f_{\phi}(z)=\lim_{n\rightarrow\infty}f_{n}(z),z\in\Delta$, we see that $f_{\phi}\in\mathcal{S}$ with $\phi(z)=S_{f_{\phi}}(z)$. In view of the normalization, we believe that $f_{\phi}$ should be unique, here the statement $f_{\phi}$ is unique means that, if there is another sequence $\{\widehat{f}_{n}(z)\}_{n=1}^{\infty}$, $\widehat{f}_{n}\in\mathcal{S}_{q}^{\infty}$ such that $\lim_{n\rightarrow\infty}\|S_{\widehat{f}_{n}}-\phi\|_{E_{2}}=0,$ and for each $z\in\Delta$, the sequence $\{\widehat{f}_{n}(z)\}_{n=1}^{\infty}$ converges, then, take $\widehat{f}_{\phi}(z)=\lim_{n\rightarrow\infty}\widehat{f}_{n}(z)$, we have $\widehat{f}_{\phi}(z)=f_{\phi}(z)$ for any $z\in\Delta$. But we have not found a proof for this statement. We leave it as the following conjecture, which appears to be nontrivial.